# Lightweight Deep Learning-Based Channel Estimation for RIS-Aided Extremely Large-Scale MIMO Systems on Resource-Limited Edge Devices

### v1

@@ IV Spatial Correlation
Spatial correlation refers to the degree of similarity between signals received or transmitted across different antenna elements, primarily influenced by their physical proximity and exposure to similar propagation paths [19]. When the RIS elements and BS antennas are closely spaced, the angular spread of the incoming and outgoing signals becomes narrow, resulting in similar fading patterns across the array. Consequently, the cascaded channel experiences a certain level of spatial correlation [20]. This section provides a detailed discussion on the spatial correlation present in the cascaded channel and explores how it can be effectively leveraged to optimize the training process and reduce computational complexity.
@@ IV-A Spatial Correlation Between RIS Elements
Let $\mathbf{R}_{r}\in\mathbb{C}^{N\times N}$ represent the correlation matrix of the RIS, which can be expressed as $\mathbf{R}_{r}=\mathbb{E}\left\{\mathbf{h}_{m}\mathbf{h}_{m}^{H}\right\}=\mathbb{E}\left\{\mathbf{a}_{r}(\phi,\theta)\mathbf{a}_{r}(\phi,\theta)^{H}\right\}$ [18]. To determine the correlation between elements, we select another RIS element $n^{\prime}$ and analyze the relationship between their respective channel responses, which can be expressed as:
@@
$\quad\cdot\left(e^{j\frac{2\pi}{\lambda}\big{(}\alpha_{n^{\prime}}\cos(\psi)\sin(\phi)d_{H}+\gamma_{n^{\prime}}\sin(\phi)d_{V}\big{)}}\right)^{H}\Bigg{]},$ (22)
@@
$\!\!\mathbf{R}_{r_{[n,n^{\prime}]}}\!\!\!\!\!=\!\!\mathbb{E}\!\left[e^{j\frac{2\pi}{\lambda}\left[(\alpha_{n}\!-\alpha_{n^{\prime}})\!\cos(\psi)\sin(\phi)d_{H}+(\gamma_{n}\!-\gamma_{n^{\prime}})\sin(\phi)d_{V}\right]}\right]\!,\!\!\!\!$ (23)
We consider the case where $\alpha_{n^{\prime}}=\alpha_{n}$, which means that the elements share the same row. This simplifies the expression for $\mathbf{R}_{r_{[n,n^{\prime}]}}$, as the term that involves $N_{h}$ disappears. The expectation can be further expressed using Equation (3) as follows:
@@
$\mathbf{R}_{r_{[n,n^{\prime}]}}=J_{0}\left({\frac{2\pi}{\lambda}||\boldsymbol{d}_{n}-\boldsymbol{d}_{n^{\prime}}||}\right),$ (26)
The correlation function $\mathbf{R}_{r_{[n,n^{\prime}]}}$ follows a zero-order Bessel function of the first kind [20]. This expression captures the correlation between two RIS elements separated by $||\boldsymbol{d}_{n}-\boldsymbol{d}_{n^{\prime}}||$. The Bessel function structure indicates that the correlation oscillates and gradually decays with increasing element separation, reflecting the spatial coherence properties of the scattered waves.
@@ IV-B Spatial Correlation Between RIS and BS
Note that equation (27) follows a structure similar to equation (21), with the key distinction being in the variance terms $[\mathbf{R}_{b}]_{m,m}$ and $[\mathbf{R}_{b}]_{m^{\prime},m^{\prime}}$. These variance terms capture the spatial correlation among antennas, which are not independent but follow the correlation function $[\mathbf{R}_{b}]_{m,m^{\prime}}=\rho^{|m-m^{\prime}|}$, where $\rho$ represents the correlation coefficient. In conclusion, the channel matrix $\mathbf{G}_{k}\in\mathbb{C}^{M\times N}$ exhibits spatial correlation, governed by the correlation structures of RIS and BS. Thus, in this paper, we leverage spatial correlation to minimize training complexity by proposing a patch-based training framework, which is discussed in detail in the following section.
@@ V Patches based Training of Channel State Information
Training machine learning models on edge devices is becoming increasingly popular, driven by the need for low-latency performance and enhanced data privacy [17]. Furthermore, it facilitates the development of robust and adaptive models that are tailored to local conditions, improving the overall performance of edge applications. However, training RIS-enabled XL-MIMO models on edge devices is challenging due to the size of the cascaded channel estimation matrices, $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$, where both $M$ and $N$ can range from hundreds to thousands [4]. This results in massive data sizes, making storing and processing such large amounts of data on resource-constrained edge devices infeasible. To address this challenge, we propose a patch-based localized training mechanism. This method leverages spatial correlation between channels to reduce computational and memory demands, enabling feasible model training on edge devices despite their limited resources.
@@ V-A Patches Based Training Framework
The patch-based training framework addresses the challenges of training large models on resource-constrained devices. In this framework, instead of processing the entire high-resolution data point $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$ at once, the proposed approach divides it into smaller, more manageable patches. Each patch is trained separately, which reduces the memory and computational requirements, making the training process more efficient. Patch-based training is beneficial for applications in edge devices, where computational and storage resources are limited. In the proposed framework, we exploit the spatial correlation present in the cascaded channel to observe that extracting a small number of representative patches from each training example is sufficient for effective model training. These extracted patches from each data point capture the essential features required for training, eliminating the need to process the entire high-dimensional matrix $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$. This is particularly relevant in the context of RIS-assisted XL-MIMO systems, where the number of RIS elements can range from hundreds to thousands. Our proposed approach significantly reduces computational complexity by focusing on smaller patches. This reduction in complexity enables resource-constrained edge devices to perform training at the edge, which would otherwise be infeasible when dealing with large-scale datasets.
As illustrated in Algorithm 1, patches are extracted from each data point in the training set $X=\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}$ along with their corresponding labels $Y={\mathbf{G}}^{(i)}_{k}$ to train the model. The inputs to the algorithm include the training set $X$, the labels $Y$, the patch dimensions $p_{x}$ and $p_{y}$, and the total number of patches total_patches to be generated. Initially, the algorithm calculates the number of patches per data point $p_{\text{dp}}$. The algorithm then iterates over the entire training set and generates $p_{\text{dp}}$ patches for each data point. For each patch, random starting coordinates $x$ and $y$ are selected within the valid range defined by the data dimensions $M$ and $N$. Subsequently, the training samples $X_{dp}$ and labels $Y_{dp}$ are sliced at the selected coordinates to extract patches of size $p_{x}\times p_{y}$ from the data point along with their corresponding labels. These patches are then appended to the lists patches_data and patches_labels. Finally, the algorithm returns patches and labels for use in model training.
Algorithm 1 Patch-Based Training Data Generation
[…]
3:  $p_{\text{dp}}=\frac{\text{total\_patches}}{\text{length}(X)}$
[…]
5:  for  $i=1$ to ${\text{length}(X)}$ do
6:     for  $j=1$ to $p_{\text{dp}}$ do
7:        $x,y\leftarrow\text{rand}(1,N-p_{x}),\text{rand}(1,M-p_{y})$
[…]
10:     end for
11:  end for
[…]
@@ V-B Efficacy of the Proposed Patch-Based Model Training
Using the entire data point $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$ for model training may not be the most efficient approach, given the inherent spatial correlation between antennas and RIS elements in the received signal at the BS, as discussed in the spatial correlation section [25]. From equations (21) and (27), it is evident that the signals received at the $m$-th and $m^{\prime}$-th antennas of the BS exhibit a significant degree of correlation. Consequently, $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$ share spatial correlation, and training the deep learning model with $\mathbf{h}_{m}$ alone is sufficient. The model can inherently learn the correlation structure of $\mathbf{h}_{m^{\prime}}$, which means that there is no need to explicitly include $\mathbf{h}_{m^{\prime}}$ in the training process.
Deep learning models, especially UNet 3+, are particularly suited for this task because they can capture spatial patterns [22]. In our proposed framework, the encoder compresses the input $\mathbf{h}_{m}$ into a lower-dimensional latent space at the bottleneck, capturing the most important features and correlations in the data. This latent representation inherently encodes the spatial correlation between $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$. The decoder then reconstructs the output from this compact representation. However, since the spatial correlation is implicitly captured in the latent space, the decoder does not need direct access to $\mathbf{h}_{m^{\prime}}$ to understand its structure. The model can generalize and learn the dependencies, making it unnecessary to include both $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$ during training. This reduces the complexity of the model while still allowing it to learn the underlying correlation pattern, ensuring that the model performs effectively even with extremely large-scale RIS-based MIMO systems.
@@ VI Computational Complexity
The computational complexity of the proposed method mainly depends on the encoder’s complexity and the decoder’s architecture. The complexity of the encoder in the proposed UNet 3+ based architecture is primarily governed by the convolutional operations and the downsampling (pooling) steps. At each encoder level, two convolutional layers are applied, followed by a max-pooling operation. Given an input feature map of dimensions $H\times W\times C^{\text{in}}$, a convolutional layer with $C^{\text{out}}$ filters and a kernel size of $K\times K$ incurs a computational complexity of $\mathcal{O}(HWC^{\text{in}}C^{\text{out}}K^{2})$. Since each level in the encoder of the proposed architecture comprises two convolutional layers, along with a $1\times 1$ convolution for dimensional alignment and a max-pooling operation, the total computational complexity at level $l$, denoted by $\mathcal{O}_{l}^{e}$, is given by:
$\mathcal{O}_{l}^{e}=\mathcal{O}(2H_{l}W_{l}C^{\text{in}}_{l}C^{\text{out}}_{l}K^{2}+2H_{l}W_{l}C^{\text{out}}_{l})$ (28)
Given that the term $2H_{l}W_{l}C^{\text{out}}_{l}$ contributes relatively little to the overall complexity, it can be considered negligible in the computational cost analysis. Furthermore, the max-pooling operation reduces the spatial dimensions by a factor of two at each level, while the number of feature channels doubles. Thus, for level $l$, the dimensions of the feature map evolve as follows:
@@
where $H_{0}\times W_{0}$ represents the spatial dimensions of the input feature map. By summing over all levels from $l=0$ to $L-1$, the total computational complexity of the encoder, represented as $\mathcal{O}^{e}$, can be expressed as:
$\mathcal{O}^{e}=\mathcal{O}\left(\sum_{l=0}^{L-1}2\frac{H_{0}W_{0}}{4^{l}}C^{in}_{l}(2C^{in}_{l})K^{2}\right),$ (30)
@@
$\approx\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L\right)$ (31)
Eq. (31) highlights that the computational cost of the encoder scales linearly with the spatial resolution $H_{0}\times W_{0}$, and quadratically with the number of output filters $\left(C_{0}^{\text{out}}\right)$. Furthermore, deeper networks result in increased computational overhead due to the growing number of feature maps. Moreover, the complexity of the bottleneck layer adopts the same convolutional structure, denoted by $\mathcal{O}^{b}$, is given by:
$\mathcal{O}^{b}=\mathcal{O}\left(H_{L}W_{L}(C^{in}_{l})^{2}K^{2}\right)\approx{H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}},$ (32)
The computational complexity of the decoder in the UNet3+ architecture follows a pattern similar to the encoder, but in reverse. At each decoder level, feature maps are upsampled/downsampled and concatenated before being processed by convolutional layers. The number of output channels decreases as we move towards the output layer, the complexity of $l$-th decoder layer $\mathcal{O}_{l}^{d}$ expressed as :
$\mathcal{O}_{l}^{d}=\mathcal{O}(H_{l}W_{l}C_{\text{concat}}C^{in}_{l}K^{2}),$ (33)
In the decoder, each level processes feature maps of size $H_{l}\times W_{l}$ with $C^{in}_{l}$ channels. Furthermore, the concatenation operation aggregates features from multiple levels, increasing the channel dimension. As a result, the total number of channels after concatenation becomes $C_{\text{concat}}=L\times C^{in}_{l}$. Consequently, the total decoder complexity $\mathcal{O}^{d}$ is given by:
@@
$\mathcal{O}^{d}=\mathcal{O}\left(H_{0}W_{0}(C_{0}^{out})^{2}K^{2}L^{2}\right),$ (35)
@@
$\mathcal{O}_{\text{total}}=$ $\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L\right)+{H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}}$
$+\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L^{2}\right),$ (36)
From (VI), it is evident that the overall computational complexity is directly proportional to the size of the input feature map $H_{0}\times W_{0}$. This implies that reducing the spatial dimensions of the input features can significantly reduce the computational cost of the proposed model.
VII Simulation
After reviewing the related literature, we identified [11] and [14] as the most relevant to our work. While these studies have made a satisfactory contribution to cascaded channel estimation, they are limited by using a small number of antennas and elements. Specifically, their approach involves processing large matrices $\tilde{\mathbf{G}}^{(i)}_{k}\in\mathbb{C}^{M\times N}$ for each data point and training on a dataset containing over 100,000 samples. Given the scale and complexity of XL-MIMO systems, where the number of antennas and RIS elements can be in the order of hundreds or even thousands, extending these approaches to such large-scale systems is not a feasible option due to their computational and memory demands. Consequently, we determined that a fair comparison between our proposed method and these works would not be feasible.
Moreover, we conducted an extensive literature review and found no existing work on training RIS-based XL-MIMO systems. However, we found relevant works mentioned in [26], [27], and [7], which can be used for comparison with our results. Therefore, in this paper, we compare our proposed method with the LS method [26], LMMSE method [27], and B-LMMSE method [7] to evaluate the significance of our results. In the B-LMMSE method, we utilize the binary reflection control strategy proposed in [7], and subsequently apply the LMMSE estimator for channel estimation. The performance of the proposed scheme is measured in terms of the normalized mean squared error (NMSE) $=\frac{E(\|G-\hat{G}_{\text{proposed}}\|_{F}^{2})}{E(\|G\|_{F}^{2})},$ where $\hat{G}_{\text{proposed}}$ and $G$ represent the channel estimated by the proposed method and the ground truth channel, respectively.
In this paper, we generated a synthetic dataset based on the system models described in Section III. The dataset was created while considering the memory constraints of edge devices. Specifically, a total of 10,000 data samples were generated, and from each sample, a representative data patch was extracted to encapsulate the underlying characteristics of the corresponding data point. As the antenna array has dimensions of $32\times 32$ and the RIS elements are arranged as $16\times 8$, the patch sizes were selected as $16\times 32$ and $32\times 32$ to align with the spatial correlation properties between the antenna and RIS elements. The dataset was split into training and validation sets, with 70% allocated for training and 30% for validation. Additionally, we created a separate testing dataset for different SNRs, with a sample size of 2,000 for each SNR value. During the training phase, the model was provided with a representative patch extracted from each data sample to reduce the input size, thereby addressing the computational constraints of resource-limited edge devices. In contrast, full-resolution samples ($\tilde{\mathbf{G}}^{(i)}_{k}\in\mathbb{C}^{M\times N}$) were used during testing to evaluate the model’s performance.
TABLE I: Simulation parameters

### v2

@@ IV Spatial Correlation
Spatial correlation refers to the degree of similarity between signals received or transmitted across different antenna elements, primarily influenced by their physical proximity and exposure to similar propagation paths [sp2]. When the RIS elements and BS antennas are closely spaced, the angular spread of the incident and reflected signals becomes narrow, leading to similar fading patterns across the arrays. Consequently, the cascaded channel experiences a certain level of spatial correlation [sp3]. This section provides a detailed discussion on the spatial correlation present in the cascaded channel and explores how it can be effectively leveraged to optimize the training process and reduce computational complexity.
@@ IV-A Spatial Correlation Between RIS Elements
where $\mathbf{R}_{r}\in\mathbb{C}^{N\times N}$ represent the correlation matrix of the RIS, which can be expressed as $\mathbf{R}_{r}=\mathbb{E}\left\{\mathbf{h}_{m}\mathbf{h}_{m}^{H}\right\}=\mathbb{E}\left\{\mathbf{a}_{r}(\phi,\theta)\mathbf{a}_{r}(\phi,\theta)^{H}\right\}$ [sp1]. To determine the correlation between elements, we select another RIS element $n^{\prime}$ and analyze the relationship between their respective channel responses, which can be expressed as:
@@
$\quad\cdot\left(e^{j\frac{2\pi}{\lambda}\big(\alpha_{n^{\prime}}\cos(\psi)\sin(\phi)d_{H}+\gamma_{n^{\prime}}\sin(\phi)d_{V}\big)}\right)^{H}\Bigg].$ (22)
@@
$\!\!\mathbf{R}_{r_{[n,n^{\prime}]}}\!\!\!\!\!=\!\!\mathbb{E}\!\left[e^{j\frac{2\pi}{\lambda}\left[(\alpha_{n}\!-\alpha_{n^{\prime}})\!\cos(\psi)\sin(\phi)d_{H}+(\gamma_{n}\!-\gamma_{n^{\prime}})\sin(\phi)d_{V}\right]}\right]\!.\!\!\!\!$ (23)
We consider the case where $\alpha_{n^{\prime}}=\alpha_{n}$, which means that the elements share the same row. This simplifies the expression for $\mathbf{R}_{r_{[n,n^{\prime}]}}$, as the term that involves $N_{h}$ disappears. The expectation can be reformulated using Equation (3) as:
@@
[…]
The correlation function $\mathbf{R}_{r_{[n,n^{\prime}]}}$ follows a zero-order Bessel function of the first kind [sp3]. This expression captures the correlation between two RIS elements separated by $\boldsymbol{d}_{n}-\boldsymbol{d}_{n^{\prime}}||$. The Bessel function shows that the correlation oscillates and gradually decays with increasing element separation, reflecting the spatial coherence properties of the scattered waves.
@@ IV-B Spatial Correlation Between RIS and BS
Note that equation (27) follows a structure similar to equation (21), with the key distinction being in the variance terms $\mathbf{R}_{b}]_{m,m}$ and $\mathbf{R}_{b}]_{m^{\prime},m^{\prime}}$. These variance terms capture the spatial correlation among antennas, which are not independent but follow the correlation function $\mathbf{R}_{b}]_{m,m^{\prime}}=\rho^{|m-m^{\prime}|}$, where $\rho$ represents the correlation coefficient. In conclusion, the channel matrix $\mathbf{G}_{i}\in\mathbb{C}^{M\times N}$ exhibits spatial correlation, governed by the correlation structures of RIS and BS. Thus, in this paper, we leverage spatial correlation to minimize training complexity by proposing a patch-based training framework, which is discussed in detail in the following section.
@@ V Patches based Training of Channel State Information
Training machine learning models on edge devices is becoming increasingly popular, driven by the need for low-latency performance and enhanced data privacy [FL]. Furthermore, it facilitates the development of robust and adaptive models that are tailored to local conditions, improving the overall performance of edge applications. However, training RIS-enabled XL-MIMO models on edge devices is challenging due to the size of the cascaded channel estimation matrices, $\tilde{\mathbf{G}}^{\text{LS}}_{i}\in\mathbb{C}^{M\times N}$, where both $M$ and $N$ can range from hundreds to thousands [XL3]. This results in massive data sizes, making storing and processing such large amounts of data on resource-constrained edge devices infeasible. To address this challenge, we propose a patch-based localized training mechanism. This method leverages spatial correlation between channels to reduce computational and memory demands, enabling feasible model training on edge devices despite their limited resources.
@@ V-A Patches Based Training Framework
The proposed patch-based training framework effectively addresses the challenges associated with training large models on resource-constrained devices. Rather than processing the entire high-resolution data point $\tilde{\mathbf{G}}^{\text{LS}}_{i}\in\mathbb{C}^{M\times N}$ at once, the framework partitions it into smaller, more manageable patches, as illustrated in Fig. 4. Training on these low-dimensional patches substantially reduces both memory and computational requirements, thereby enhancing overall training efficiency. This patch-based approach is particularly advantageous for edge devices, where computational and storage resources are limited. In addition, by exploiting the spatial correlation inherent in the cascaded channel, it is observed that a small number of representative patches per training example suffices for effective model learning, as discussed in Section IV. These patches capture the essential features required for training, eliminating the need to process the full high-dimensional matrix $\tilde{\mathbf{G}}^{\text{LS}}_{i}\in\mathbb{C}^{M\times N}$. This is particularly beneficial in the context of RIS-assisted XL-MIMO systems, where the number of RIS elements can range from hundreds to thousands. Our proposed approach significantly reduces computational complexity by processing only a few small-dimensional patches, rather than processing the entire high-dimensional input. This reduction in complexity enables resource-constrained edge devices to perform training at the edge, which would otherwise be infeasible when dealing with high-dimensional datasets.
As illustrated in Fig. 4, $\tau$ represents the number of data points in $\tilde{\mathbf{G}}^{\text{LS}}$. From each data point $\tilde{\mathbf{G}}^{\text{LS}}_{i}$ in the training set, $\mathcal{P}$ patches of size $\mathcal{P}_{x}\times\mathcal{P}_{y}$ are extracted along with their corresponding labels $\mathbf{G}_{i}$ to train the model, where $\mathcal{P}_{x}$ and $\mathcal{P}_{y}$ denote the patch dimensions along the $x$- and $y$-axes, respectively. We observe a trade-off between the number of patches $\mathcal{P}$ extracted from each data point, their dimensions $\mathcal{P}_{x}$ and $\mathcal{P}_{y}$, and the available computational resources, including storage and processing capabilities. Nevertheless, increasing these parameters generally improves estimation accuracy, as elaborated in the simulation results section. However, they must be carefully chosen to satisfy the resource constraints of edge devices. For each patch, random starting coordinates $x$ and $y$ are selected within the valid range defined by the data dimensions $M$ and $N$, ensuring that the extracted patches do not overlap. This random but controlled selection strategy enhances the model’s ability to learn spatially diverse features from different regions of the input data. After the patch generation process, a total of $\mathcal{P}_{\text{Total}}=\mathcal{P}\times\text{(Total Training Data)}$ patches are generated and subsequently utilized for training. During the offline training phase, these patches and their corresponding labels are fed into the proposed architecture to optimize its parameters. Once the training is complete, the trained model is deployed for online inference. In contrast to the training stage, where the data are divided into smaller patches, the proposed architecture is designed to directly process the entire input channel matrix ${\mathbf{G}}^{\text{LS}}_{k}\in\mathbb{C}^{M\times N}$ during inference, producing the estimated channel $\tilde{\mathbf{G}}_{i}\in\mathbb{C}^{M\times N}$ without the need for patch extraction. This design choice significantly reduces inference time while maintaining high estimation accuracy.
Figure 4: The training and inference stages of the proposed framework use matrices of dimension $\mathbf{R}$, where the last dimension represents the real and imaginary parts of the channel.
@@ V-B Efficacy of the Proposed Patch-Based Model Training
Training the model using high-dimensional data points $\tilde{\mathbf{G}}^{\text{LS}}_{i}\in\mathbb{C}^{M\times N}$ may not be the most efficient approach, given the inherent spatial correlation between antennas and RIS elements in the received signal at the BS, as discussed in the section III. From equations (21) and (27), it is evident that the signals received at the $m$-th and $m^{\prime}$-th antennas of the BS exhibit a significant degree of correlation. Consequently, $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$ share spatial correlation, and training the deep learning model with $\mathbf{h}_{m}$ alone is sufficient. The model can inherently learn the correlation structure of $\mathbf{h}_{m^{\prime}}$, which means that there is no need to explicitly include $\mathbf{h}_{m^{\prime}}$ in the training process. In our proposed framework, the encoder compresses the input $\mathbf{h}_{m}$ into a lower-dimensional latent space at the bottleneck, capturing the most important features and correlations in the data. This latent representation inherently encodes the spatial correlation between $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$. The decoder then reconstructs the output from this compact representation. However, since the spatial correlation is implicitly captured in the latent space, the decoder does not need direct access to $\mathbf{h}_{m^{\prime}}$ to understand its structure. The model can generalize and learn the dependencies, making it unnecessary to include both $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$ during training. This reduces model complexity while preserving its ability to learn the underlying correlation patterns of high-dimensional data points $\tilde{\mathbf{G}}^{\text{LS}}_{i}\in\mathbb{C}^{M\times N}$, ensuring effective performance in extremely large-scale RIS-based MIMO systems.
@@ VI Computational Complexity
The computational complexity of the proposed method primarily depends on the encoder and decoder architectures. In the encoder, complexity is mainly driven by convolutional operations and downsampling, where each level applies two convolutions followed by max pooling. Given an input feature map of dimensions $H\times W\times C^{\text{in}}$, a convolutional layer with $C^{\text{out}}$ filters and a kernel size of $K\times K$ incurs a computational complexity of $\mathcal{O}(HWC^{\text{in}}C^{\text{out}}K^{2})$. Since each level in the encoder of the proposed architecture comprises two convolutional layers, along with a $1\times 1$ convolution for dimensional alignment and a max-pooling operation, the total computational complexity at level $l$, denoted by $\mathcal{O}_{l}^{e}$, is given by:
$\mathcal{O}_{l}^{e}=\mathcal{O}(2H_{l}W_{l}C^{\text{in}}_{l}C^{\text{out}}_{l}K^{2}+2H_{l}W_{l}C^{\text{out}}_{l}),$ (28)
Since $2H_{l}W_{l}C^{\text{out}}_{l}$ has negligible impact, it is omitted. The max-pooling operation halves the spatial dimensions at each level while doubling the feature channels, leading to the following feature map dimensions at level $l$:
@@
where $H_{0}\times W_{0}$ represents the spatial dimensions of the input feature map. By summing over all levels from $l=0$ to L1L-1, the total computational complexity of the encoder, represented as $\mathcal{O}^{e}$, can be expressed as:
$\mathcal{O}^{e}=\mathcal{O}\left(\sum_{l=0}^{L-1}2\frac{H_{0}W_{0}}{4^{l}}C^{in}_{l}(2C^{in}_{l})K^{2}\right).$ (30)
@@
$\approx\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L\right).$ (31)
TABLE I: Computational Complexity of Different Estimation Algorithms
Algorithms CDRN SSL Proposed
FLOPs $4.54\times 10^{12}$ $1.55\times 10^{11}$ $6.73\times 10^{7}$
Eq. (31) shows that the encoder’s computational cost scales linearly with the spatial resolution $H_{0}\times W_{0}$ and quadratically with the number of output filters $C_{0}^{\text{out}}$. Moreover, the complexity of the bottleneck layer adopts the same convolutional structure, denoted by $\mathcal{O}^{b}$, is given by:
$\mathcal{O}^{b}=\mathcal{O}\left(H_{L}W_{L}(C^{in}_{l})^{2}K^{2}\right)\approx{H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}}.$ (32)
The decoder’s computational complexity mirrors the encoder in reverse, with feature maps up/downsampled, concatenated, and convolved, while output channels decrease toward the final layer. The complexity of the $l$-th decoder layer $\mathcal{O}_{l}^{d}$ is given by:
$\mathcal{O}_{l}^{d}=\mathcal{O}(H_{l}W_{l}C_{\text{concat}}C^{in}_{l}K^{2}).$ (33)
In the decoder, each level processes feature maps of size $H_{l}\times W_{l}$ with $C^{\text{in}}_{l}$ channels. Concatenation aggregates features from multiple levels, increasing the channel dimension to $C_{\text{concat}}=L\times C^{\text{in}}_{l}$. The total decoder complexity is then given by $\mathcal{O}^{d}$:
@@
$\mathcal{O}^{d}=\mathcal{O}\left(H_{0}W_{0}(C_{0}^{out})^{2}K^{2}L^{2}\right).$ (35)
@@
$\mathcal{O}_{\text{total}}=\$$\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L\right)+{H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}}$
$+\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L^{2}\right).$ (36)
From (VI), it is evident that the overall computational complexity scales directly with the size of the input feature map $H_{0}\times W_{0}$, indicating that leveraging a patch-based training approach can substantially reduce the computational cost of the proposed model. Moreover, Table I shows that the benchmark schemes CDRN [DL1] and SSL [DL4] require significantly more floating-point operations (FLOPs) per input sample during training compared to our proposed patch-based framework.
TABLE II: Simulation parameters